OpenStudy (anonymous):

Solve the following using Cramer's Rule: y = 2x - 13 y = -x + 5

OpenStudy (anonymous):

Ok, this is probably the hard way to solve this system, but here we go. First rearrange the given system to a setup friendlier for changing into a matrix. After a little work, we will get\[2x-y=13\]\[x+y=5\]

OpenStudy (anonymous):

You with me so far?

OpenStudy (anonymous):

Now, we need to change the system into a matrix. The LHS becomes\[A=\left[\begin{matrix}2 &-1\\ 1 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

We will substitute the RHS vector (the two numbers right of the equal sign in for the x column and the y column to make A_x and A_y respectively. Those matrices are\[A_x=\left[\begin{matrix}13 & -1\\ 5 & 1\end{matrix}\right]\]\[A_y=\left[\begin{matrix}2 & 13 \\ 1 & 5\end{matrix}\right]\]

OpenStudy (anonymous):

Now we apply Cramer's rule. To solve the system we substitute, take some determinants, and get the solution (x,y) for the system.\[x=\frac{detA_x}{detA}=\frac{18}{3}=6\]\[y=\frac{detA_y}{detA}=\frac{-3}{3}=-1\]\[(x,y)=(6,-1)\]

OpenStudy (anonymous):

We can check the answer by solving using the elimination method, which gives the identical solution.

OpenStudy (anonymous):

Hope you understand this...

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No sweat.